OpenStudy (anonymous):

. What is the slope of a line perpendicular to the line with equation y = 5x 2? (Points : 4) - 5 5 -1/5 -2

OpenStudy (freckles):

if y=5x^2 then we don't have a line. is y=5x(2) 2(5)=10 so y=10x

OpenStudy (freckles):

so if we have the line y=10x what is the slope of y=10x?

OpenStudy (freckles):

the line that is perpendicular to it will have opposite recipocal slope

OpenStudy (anonymous):

0_0

OpenStudy (freckles):

Can you tell me what the face means?

OpenStudy (freckles):

I know it is sad or something but let me know if you don't understand something.

OpenStudy (freckles):

For some reason I don't think you really meant to write y=5x 2

OpenStudy (anonymous):

0_0 means im confused XD its 0_0 or 0.0

OpenStudy (anonymous):

yea i did cause thats what the question is i copyed and pasted XD

OpenStudy (freckles):

I guess there isn't anything else you forgot to copy?

OpenStudy (freckles):

like a plus sign or minus sign

OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 5x – 2? (Points : 4)

OpenStudy (freckles):

y=5x 2 usually means y=5x^2 but that wouldn't make since because they are calling y=5x 2 a line not a quadratic y=5x 2 could also be interpreted as y=5x(2)=5(2)x=10x

OpenStudy (anonymous):

exact copy

OpenStudy (anonymous):

lemme screen shot it

OpenStudy (freckles):

no you copied the sign in between

OpenStudy (freckles):

i guess you don't understand what i was saying you wrote y=5x 2 earlier now you are writing y=5x-2

OpenStudy (anonymous):

oh XD

OpenStudy (freckles):

What is the slope of y=5x-2?

OpenStudy (anonymous):

this isnt even my stuff 0_0

OpenStudy (freckles):

the slope of y=mx+b is m

OpenStudy (freckles):

so the slope of y=5x-2 is?

OpenStudy (freckles):

Can I talk to the person who is studying this stuff?

OpenStudy (anonymous):

no hes gone for a while

OpenStudy (freckles):

well anyways the slope of y=mx+b is m (m not equal to 0) any line that is perpendicular to y=mx+b will have slope -1/m -- lines not included above are vertical lines and horizontal lines... lines that are perpendicular to horionztal lines are vertial lines lines that are perpendiculart to vertical lines are horizontal lines